OpenStudy (anonymous):

Hi everyone, my teacher just gave us this problem and nobody was able to solve it! It's a 2-dimension kinematics problem. A projectile is fired to the East with an initial velocity of 35m/s and lands 184 m away. What angle was the projectile fired at with respect to the x-axis? Good luck! :o)~

OpenStudy (anonymous):

I can tell you that a double angle formula is supposed to be used but after 5 hours trying every combination of trig function from my calc book, I cannot solve this problem!

OpenStudy (anonymous):

i know Vx=35cos(theta) and Vy=35sin(theta)

hartnn (hartnn):

Horizontal range R = u^2 sin 2theta/g u = initial velocity

OpenStudy (anonymous):

deltaX= 184m also Xi = 0 Xf= 184

OpenStudy (anonymous):

so 35^2sin2theta/-9.8 is that right?

hartnn (hartnn):

=184, yes

OpenStudy (anonymous):

so i can solve for theta from that?

hartnn (hartnn):

nopes, something is missing...

OpenStudy (anonymous):

sin(2theta)=-1.472

hartnn (hartnn):

which is not possible

OpenStudy (anonymous):

thats what i thought...its like not enough information but she told everyone she wanted to see how sharp we were...i was like whatever lady!

OpenStudy (anonymous):

so should i just give up? she claims that we should have used theta/2 as angle under some radical...very confused

hartnn (hartnn):

idk bout that but i will give one more shot, by finding the time 't'

OpenStudy (anonymous):

i even tried spliting the prblem into half by just focusing on the max height and dividing mt range by 2 184/2=92 but that got me nowhere either

OpenStudy (anonymous):

I even put 35sintheta for my Viy and 35costheta for my Vix and then equated both of those with time but it still didnt work

hartnn (hartnn):

i got imaginary time :O i had to use a= +9.8 to get real time, but that wouldn't be correct

OpenStudy (anonymous):

wow...maybe she is crazy

OpenStudy (anonymous):

here is a question...even if we made our own hypothetical problem where the angle was 45 degrees, by symmetry, we would know the the final velocity and angle would also be the same, but without either know the time or the distance, we still can't create our own problem in hopes of relating the two can we?

hartnn (hartnn):

with initial vel as 35 m/s the maximum dist. the projectile can travel (taking theta=45) will be just 125 m it would not go 184 m far distance and time are related as s=ut+0.5at^2 if thats what u were asking

OpenStudy (anonymous):

well thanks for giving it a try!

OpenStudy (badhi):

Maybe it was shot from an elevated place ;)

OpenStudy (anonymous):

Hi badhi...how would that have made a difference?

OpenStudy (anonymous):

ok...lets say hypothetically, 28m cliff...you could solve it then?

OpenStudy (badhi):

 If we write equations, for many \((\theta,x)\) values we can get the required distance - 184. Which will make the question meaningless. But for me it seems to be the only explanation.